OpenStudy (anonymous):

Find the distance between the points (2.7, 5.1) and (3, 4.7). 0.05 0.5 5

OpenStudy (anonymous):

What?

OpenStudy (littlebird):

I just took a picture since I noticed the mistake I made.

OpenStudy (anonymous):

I can't get pictures

OpenStudy (ja1):

\[Distance = \sqrt{(X _{2} - X_{1})^{2} + (Y_{2} - Y_{1})^{2}}\]

OpenStudy (ja1):

Plug 'em in

OpenStudy (anonymous):

I know that part. But the answer I got from that isn't one of the choices.

OpenStudy (littlebird):

It worked for me

OpenStudy (anonymous):

What answer did you get? I got something like 0.7071067... Yeah...

OpenStudy (ja1):

Ok let's work it out together, first stack them.

OpenStudy (ja1):

Or you could do that....

OpenStudy (littlebird):

Should I let you figure it out or put it back..?

OpenStudy (littlebird):

(2.7-3)^2=.09 (5.1-4.7)^2=.16 .09+.16=.25 sqrt(.25)=.5 Sorry! I'm always worrying about breaking the websites rules about answering questions. :P here it is again

OpenStudy (anonymous):

Thank you:) I appreciate it!

OpenStudy (ja1):

@Littlebird Well you are not breaking it since you showed your work but it's always best to guide the user and never give them the answer, as the old saying goes, if you give a man a fish he will eat for a night, but if you teach him to fish he will eat every day ;)